OpenStudy (elisaneedshelp):

The hour hand is 3 on the Springfield city hall clock. This clock has a minute hand of 10 feet. If the shortest distance from the end of the minute hand to the hour hand is 6.5 feet, what time is it? (Two possibilities) Round to the nearest minute.

OpenStudy (elisaneedshelp):

@nincompoop can you help me with this one too?

OpenStudy (elisaneedshelp):

@campbell_st @Luigi0210 @whpalmer4 please

OpenStudy (elisaneedshelp):

Please i have very limited time and i need help

OpenStudy (campbell_st):

what grade..?

OpenStudy (elisaneedshelp):

Im in 10th grade please help me

OpenStudy (elisaneedshelp):

This problem has something to do with trigonometry but idk how to do it myself

OpenStudy (campbell_st):

well this is what I'd say  find the size of angle x sin(x) = 6.5/10 then get an angle...

OpenStudy (campbell_st):

when you know the angle use angle/360 * 60 will give the the number of minutes.... so the 2 answers will be 15 - minutes xx 15 + minutes yy to get the time... it will be 3: xx and 3:yy

OpenStudy (campbell_st):

so what did you get for the angle..?

OpenStudy (elisaneedshelp):

i dont think i got it right at all....0.65...i guess i didnt put it in correctly

OpenStudy (campbell_st):

you need to find arc sin or inverse sin to get the angle. \[x = \sin^{-1}(\frac{6.5}{10})\]

OpenStudy (elisaneedshelp):

40.54

OpenStudy (elisaneedshelp):

40.54160187

OpenStudy (campbell_st):

great how many minutes is that use 40.54/360 * 60 =

OpenStudy (elisaneedshelp):

0.001876852

OpenStudy (campbell_st):

hold on you need \[(40.54 \times 60)\div 360 = \]

OpenStudy (elisaneedshelp):

6.756666667

OpenStudy (campbell_st):

great so the minute hand is 6.76 minutes above the 3 on the clock so rounding is 3:(15 - 7) oclock and after 15 its 3:(15 + 7) oclock

OpenStudy (elisaneedshelp):

3:08 and 3:22

OpenStudy (campbell_st):

that would be my answer

OpenStudy (elisaneedshelp):

yay thank you and could you help me with another one? An airplane is on a 36 degree approach to the airport.If the altimeter reads 25,000feet, how long is the remaining flight path (the diagonal) in miles? Round to the 1000th of a mile.

OpenStudy (campbell_st):

you could have used the rule of sines for that... not sure if you've done it...

OpenStudy (campbell_st):

post the question in the normal way rather than adding... as someone else will help...

OpenStudy (campbell_st):

I'm not confident on the solution just rethinking it... perhaps law of sines would have been better

OpenStudy (campbell_st):

perhaps 3:10 and 3:20 may be closer

OpenStudy (elisaneedshelp):

then we would have to change how we worked it out...

OpenStudy (campbell_st):

well my question is... have you done the law of sine... or the sine rule..?

OpenStudy (elisaneedshelp):

Nope and anyway i cant spend too much time on one problem cuz i only have two hours anyway before i have to get ready for school :/

OpenStudy (campbell_st):

ok... so I'd leave it at that

OpenStudy (elisaneedshelp):

yep